OpenStudy (anonymous):

PLEASE HELP, URGENT. Write an equation of a line passing through (-5, 9) and is perpendicular to 2x+3y=4.

OpenStudy (anonymous):

since the line is perpendicular so the product of the slopes of the two lines will be -1 slope of the given line is m=-2/3 slope of requirede line must be \[\Large slope=\frac{-1}{m}=\frac{-1}{\frac{-2}{3}}=\frac{3}{2}\] now use point slope form y-y0=m(x-x0) where m=3/2 x0=-5 yo=9 put these values in the above equation to get the equation of line

OpenStudy (anonymous):

umm.... I'm not sure what this is supposed to mean: \[\Large slope=\frac{-1}{m}=\frac{-1}{\frac{-2}{3}}=\frac{3}{2}\]

OpenStudy (anonymous):

since it is given that lines are perpendicular for the perpendicular lines with slopes m1 and m2 following must be satisfied . \[\Large m_{1}m_{2}=-1\] it means if the lines are perpendicular and you know the slope of one line line the slope of the other line will be the -1 of the reciprocal of the perdicular line to it \[\Large m_{2}=\frac{-1}{m_{1}}\] since we already know the slope of one line \[\Large m_{1}=\frac{-2}{3}\] so slope of the line perpendicular to it can be found easily . \[\Large m_{2}=\frac{-1}{m_{1}}=\frac{-1}{\frac{-2}{3}}=\frac{3}{2}\] this is slope of the required line .

OpenStudy (anonymous):

let me know if don't follow this.

OpenStudy (anonymous):

OH NVM, the equations just weren't loading, OKAY, I got it now!

OpenStudy (anonymous):

THANK YOU!